OpenStudy (anonymous):

Which type of graph will the function f(x) = e-x give? elliptical exponential parabolic hyperbolic

OpenStudy (zale101):

Figure the y intercept and x intercept and then connect the lines.

OpenStudy (zale101):

f(x)=e-x can be f(x)=-x+e hence the y=mx+b form

OpenStudy (anonymous):

that still didnt help !!!

OpenStudy (zale101):

Are you sure the function is f(x) = e-x?

OpenStudy (zale101):

because that is a line

OpenStudy (zale101):